Suppose that 32% students are addicted to chocolate. If 10 students are randomly selected, what is the probability that at least one of them is addicted to chocolate? Round your answer to five decimal places.
Answer by stanbon(57307)

Suppose that 32% students are addicted to chocolate. If 10 students are randomly selected, what is the probability that at least one of them is addicted to chocolate? Round your answer to five decimal places.
---
P(at least one addicted) = 1 - P(none addicted)
= 1 - (0.68)^10
---
= 1 - 0.021139
---
= 0.9787
=============
Cheers,
Stan H.
=============